Let f be the bilinear form on R2 defined by

f[(x1, x2), (y1, y2)]= 3x1y1-2x1y2+4x2y1-x2y2

then find the matrix A of f in the basis {u1(1, 1)=u2(1, 2)}


1
Expert's answer
2021-01-10T16:33:49-0500

"f: \Reals^2*\Reals^2\to\Reals"is given by"f(\t(\\begin{matrix}\n x_1 \\\\\n x_2\n\\end{matrix}),\n(\\begin{matrix}\n y_1 \\\\\n y_2\n\\end{matrix}))=3x_1y_1-2x_1y_2+4x_2y_1-x_2y_2"

Let us write the matrix of f in the standard basis.

"f(e_1,e_1)=3,\space f(e_1,e_2)=-2, \space f(e_2,e_1)=4, \space f(e_2,e_2)=-1"

hence the matrix in the standard basis is

"A=( \t\\begin{matrix}\n 3 & -2 \\\\\n 4 & -1\n\\end{matrix})"

Now for "B = \lbrace b_1, . . . , b_n \rbrace" and "U = \lbrace u_1, . . . , u_n\rbrace" are two bases for "V" . We may write one basis in terms of the other: "u_i= \sum_{j=1}^{n}\lambda_{j,i}b_j." Where coefficients "\lambda_{j,i}" form the transition matrix "M" from "B" to "C". To found matrix "A" in new basis "\lbrack A\rbrack_u" used change of basis formula : "\lbrack A\rbrack_u=M^T \lbrack A\rbrack_b M"

In this case we want to write this matrix i basis

"\lparen \begin{matrix}
 1\\
 1
\end{matrix})" ,"(\begin{matrix}
 1\\
 2
\end{matrix})"

The transition matrix is :

"M=( \begin{matrix}
 1 & 1 \\
 1 & 2
\end{matrix})"

it’s transpose is the same. The matrix of f in the new basis is

"\\lbrack A\\rbrack_u=( \\begin{matrix}\n 1 & 1 \\\\\n 1 & 2\n\\end{matrix}) ( \t\\begin{matrix}\n 3 & -2 \\\\\n 4 & -1\n\\end{matrix}) ( \\begin{matrix}\n 1 & 1 \\\\\n 1 & 2\n\\end{matrix})= ( \\begin{matrix}\n 4 & 1 \\\\\n 7 & 3\n\\end{matrix})"